simplify: (5^4/5 * 5^4/5)^-10
1
Expert's answer
2012-12-21T05:04:16-0500
simplify: (5^4/5 *5^4/5)^-10
(a^b)^c = a^(b+c)
1/5 = 5^(-1)
(5^4/5 * 5^4/5)^-10 = (5^(4-1) * 5^(4-1))^-10 = ( 5^3 * 5^3)^-10 =(5^6)^-10 = 5^(6-10) = 5^-4